An investor invested a total a total of $1,100 in two mutual funds.One fund earned a 5% profit while the other earned a 2% profit. If the investor's total profit was $46, how much was invested in each mutual fund?
Answer by mananth(16946)

Part I 5.00% per annum ------------- Amount invested =x
Part II 2.00% per annum ------------ Amount invested = y
1100
Interest----- 46.00

Part I 5.00% per annum ---x
Part II 2.00% per annum ---y
Total investment
x + 1 y= 1100 -------------1
Interest on both investments
5.00% x + 2.00% y= 46
Multiply by 100
5 x + 2 y= 4600.00 --------2
Multiply (1) by -5
we get
-5 x -5 y= -5500.00
Add this to (2)
0 x -3 y= -900
divide by -3
y = 300
Part I 5.00% $ 800
Part II 2.00% $ 300

CHECK
800 --------- 5.00% ------- 40.00
300 ------------- 2.00% ------- 6.00
Total -------------------- 46.00

m.ananth@hotmail.ca